# Performance of a coarsely pixelated LAPPD photosensor for the SoLID light gas Cherenkov detector

### v2

@@ 7 Summary
We reported on the performance study of a coarse-pixelated LAPPD using a telescopic Cherenkov device in an open high-rate environment at Jefferson Lab. A pixelated Gen-II LAPPD was tested with two different radiators, a light gas CO and a heavy gas CF that will be used in the SoLID experiment. The coarse-pixelated Gen-II LAPPD device was able to operate in such an environment for both gas radiators, detect the Cherenkov signals, and separate single-electron events and pair production events. With this LAPPD a typical Cherenkov disk was observed in our TCD, with a smaller diameter for CO radiator and a larger diameter for CF radiator. A larger number of photoelectrons was generated from the signal sum in CF compared to CO case as expected. GEANT 4 simulations were performed to confirm the results, showing a good agreement with the data. Our experiment confirms that for SoLID the pixelated Gen-II LAPPD could handle the high background rate environment for the more stringent requirement of the SoLID PVDIS experiment.
@@ Acknowledgement
Argonne National Laboratory’s work was supported by the US Department of Energy, Office of Science, Office of Nuclear Physics, under contract DE-AC02-06CH11357. Part of the work was supported by Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, provided by the Director, Office of Science, of the US Department of Energy under contract DE-AC02-06CH11357. The work at Temple University is supported by the US Department of Energy award DE-SC0016577. Incom’s work is supported by the U. S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Offices of High Energy Physics and Nuclear Physics under DOE contracts: DE-SC0015267, DE-SC0017929, DE-SC0018778, and DE-SC0019821. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177.

### v4

@@ 7 Summary
We report on the performance study of a coarse-pixelated LAPPD using a telescopic Cherenkov device in an open high-rate environment at Jefferson Lab. A pixelated Gen-II LAPPD was tested with two different radiators, a light gas CO2 and a heavy gas C4F8, which are planned to be used in the SoLID experiment. We demonstrate that the coarse-pixelated Gen-II LAPPD device can operate in a high-rate environment, consistent with the expected rates during the SoLID and SIDIS experiments. Furthermore, the pixelated information is critical to selecting a clean sample of the Cherenkov events from the random coincidences background. The LAPPD’s response to Cherenkov photons is studied with two gas radiators, namely CO2 and C4F8. As expected, more photoelectrons were generated from the Cherenkov process in C4F8 compared to those in CO2. GEANT4 simulations were conducted to crosscheck the results, showing a good agreement with the data. Our experiment confirms that the pixelated Gen-II LAPPD is capable of handling the high-rate environment expected for future SoLID experiments.
@@ Acknowledgement
Argonne National Laboratory’s work was supported by the US Department of Energy, Office of Science, Office of Nuclear Physics, under contract DE-AC02-06CH11357. The work at Temple University is supported by the US Department of Energy award DE-SC0016577. Incom’s work is supported by the U. S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Offices of High Energy Physics and Nuclear Physics under DOE contracts: DE-SC0015267, DE-SC0017929, DE-SC0018778, and DE-SC0019821. The work at Duke University is supported by the US Department of Energy under contract DE-FG02-03ER41231. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177.